# Probing the low-energy particle content of blazar jets through MeV observations

### v1

@@ 1 Introduction
Our aim is to explore the potential signatures of the electron thermal bump in the emission of flat-spectrum radio quasars (FSRQ). These are powerful blazars characterized by a dominant $\gamma$-ray component, probably produced through the inverse Compton scattering (IC) of ambient radiation by relativistic leptons in the jet (e.g. Sikora et al. 1994; Ghisellini & Tavecchio 2009). A back-of-the-envelope calculation suggests that, for typical parameters, electrons belonging to the thermal component emit in the MeV band, which corresponds to the maximum of the high-energy component of FSRQ (e.g. Sikora et al. 2002; Ghisellini et al. 2017; Marcotulli et al. 2022). For several of these sources the observed flux in the MeV is within the reach of the upcoming Compton Spectrometer and Imager (COSI) satellite (Tomsick & COSI Collaboration 2022), giving the opportunity to test our scenario in the near future.
@@ 2 The model
$Q(\gamma)=\begin{cases}K_{e}Q_{\rm th}(\gamma)&{\rm if}\quad\gamma<\gamma_{\rm nth}\\ K_{e}Q_{\rm th}(\gamma_{\rm nth})\left(\frac{\gamma}{\gamma_{\rm nth}}\right)^{-p}e^{-\gamma/\gamma_{\rm c}}&{\rm if}\quad\gamma>\gamma_{\rm nth}\end{cases}$ (1)
@@
$Q_{\rm th}(\gamma)=\frac{\gamma^{2}}{\gamma_{\rm th}K_{2}(1/\gamma_{\rm th})}e^{-\frac{\gamma}{\gamma_{\rm th}}},$ (2)
with $K_{2}$ the modified Bessel function of the second kind. The parameter $\gamma_{\rm th}$ plays the role of an effective temperature of the quasi-thermal distribution, while $\gamma_{\rm nth}$ in Eq.1 is the minimum Lorentz factor of the non-thermal tail, described as a power law with slope $p$ with cut-off at $\gamma_{\rm c}$ (with $\gamma_{\rm c}\gg\gamma_{\rm nth}$). These “injected” particles then cool and radiate.
@@
where $V$ is the volume of the emitting region and $m_{e}$ the electron rest mass. It is useful to introduce the fraction of energy contained in the non-thermal tail with respect to the total one:
@@
Figure 1: Electron energy distribution in the emission region calculated at different times (given in units of the light crossing time, $R/c$) with the parameters used for model A. The equilibrium distribution comprises a cooled low energy tail $N(\gamma)\propto\gamma^{-2}$, the rapidly decreasing high-energy tail of the thermal Maxwellian (which has the peak at $\gamma\simeq 100$) and the high-energy non-thermal power law with slope $p+1$ (with $p=2.3$ is the slope of the injected power law) up to the cut-off at $\gamma_{c}$.
@@ 2.1 Fixing parameters
Since we intend to explore of the effects of the complex EED on the spectral energy distribution (SED) of FSRQ, we assume benchmark values for the physical parameters of the blazar emission region inspired by modeling of FSRQ (e.g. Tavecchio et al. 2000; Ghisellini et al. 2010). We assume $R=5\times 10^{16}$ cm, $B=1.6$ G, $\Gamma=20$, $\theta_{\rm v}=3.7$ deg (which combine to give a Doppler factor $D=15$). For the external radiation we assume that the dominant radiation field is that associated to the BLR, approximated (in the source rest frame) as a black body peaking at $\nu_{\rm ext}=2\times 10^{15}$ Hz (see the discussion in Tavecchio & Ghisellini 2008) with energy density $U_{\rm ext}=2.5\times 10^{-2}$ erg cm-3 (e.g. Ghisellini & Tavecchio 2008). We further assume that the source is located at the typical redshift $z=2$.
@@
A possible scenario for the origin of the shock at which electrons are heated and subsequently accelerated to high-energies is that invoking the interaction of different portions of the flow characterized by different speed (internal shocks, e.g. Spada et al. 2001). In this case we expect a mildly relativistic shock, with $\Gamma_{\rm u}\sim 1.5-2$. These shocks can efficiently accelerate particles only for low magnetization of the plasma (Sironi et al. 2015b). Indeed, FSRQ jets at the distance where the observed radiation is produced is expected to have a small magnetization (e.g. Sikora et al. 2005; Celotti & Ghisellini 2008). In these conditions we can rely on the results of Crumley et al. (2019), that report the analysis of PIC simulations for subluminal high-Mach number, mildly relativistic ($\Gamma_{\rm u}=1.7$) weakly magnetized jets. The simulations show the development of a thermal component with average Lorentz factor around $\langle\gamma\rangle\approx 300$, and of a non-thermal power law with slope 2.2-2.3 containing a small fraction of the shock energy, $\epsilon_{e}\approx 10^{-3}$. However this last result could depend on the detailed configuration (i.e. inclination of the magnetic field and the magnetization). PIC simulations of weakly magnetized, relativistic ($\Gamma_{\rm u}>3$) shocks show that the fraction of energy conveyed to non-thermal electrons can be much larger than that estimated in the mildly relativistic case, of the order of $\epsilon_{e}\approx 10^{-1}$ (Sironi & Spitkovsky 2011). These conditions could apply to FSRQ if the acceleration occurs at a stationary oblique recollimation shock (e.g. Bodo & Tavecchio 2018; Zech & Lemoine 2021).
@@ 3 Results
Table 1: Parameters of the models reported in Fig. 2. The last row reports $\epsilon_{e}$ derived from the input parameters with Eq. 7 and $\Gamma_{\rm u}=1.5$. The injected electron luminosity $L_{e}$ is in units of $10^{44}$ erg s-1.
@@ 4 Discussion
In this paper we have studied the case of the most powerful FSRQs, for which the thermal component naturally falls in the MeV band. For other types of blazars the situation is less straightforward. Less powerful FSRQ display the EC peak at higher energies (100 MeV-1 GeV e.g. Marcotulli et al. 2022). If related to the thermal bump, such high peak energies would imply correspondingly high electron temperature, possibly related to larger dissipation parameter $\epsilon_{e}$. We remark that in this case, the high-energy part of the EC component would be affected by KN effects that smooth the spectrum, making difficult to identify the spectral structure (this effect is already visible in the D model in Fig. 2). At even lower power, blazars of the BL Lac type (LSP and HSP), display peaks above 10 GeV, which require electrons with $\gamma\gtrsim 10^{4}$ (e.g., Tavecchio et al. 2010), incompatible with thermal components from mildly relativistic shocks. In this case the thermal bump would instead appear in the low energy part of the IC component. We however remark that these sources produce high-energy radiation mainly through SSC, resulting in quite smooth spectra even for prominent thermal bumps (as also visible in Fig. 2).

### v2

@@ 1 Introduction
Our aim is to explore the potential signatures of the electron thermal bump in the emission of flat-spectrum radio quasars (FSRQ). These are powerful blazars characterized by a dominant $\gamma$-ray component, probably produced through the inverse Compton scattering (IC) of ambient radiation by relativistic leptons in the jet (e.g. Sikora et al. 1994; Ghisellini & Tavecchio 2009). A back-of-the-envelope calculation suggests that, for typical parameters, electrons belonging to the thermal component emit in the MeV band, which corresponds to the maximum of the high-energy component of FSRQ (e.g. Sikora et al. 2002; Ghisellini et al. 2017; Marcotulli et al. 2022). [^1] Baring et al. (2017) performed a similar study, but they assumed a cold thermal component, whose emission peaks in the soft X-ray band. For several of these sources the observed flux in the MeV is within the reach of the upcoming Compton Spectrometer and Imager (COSI) satellite (Tomsick & COSI Collaboration 2022), giving the opportunity to test our scenario in the near future.
@@ 2 The model
$Q(\gamma)=\begin{cases}K_{e}Q_{\rm th}(\gamma)e^{-\gamma_{\rm nth}/\gamma_{\rm c}}&{\rm if}\quad\gamma<\gamma_{\rm nth}\\ K_{e}Q_{\rm th}(\gamma_{\rm nth})\left(\frac{\gamma}{\gamma_{\rm nth}}\right)^{-p}e^{-\gamma/\gamma_{\rm c}}&{\rm if}\quad\gamma>\gamma_{\rm nth}\end{cases}$ (1)
@@
$Q_{\rm th}(\gamma)=\frac{\beta\gamma^{2}}{\gamma_{\rm th}K_{2}(1/\gamma_{\rm th})}e^{-\frac{\gamma}{\gamma_{\rm th}}},$ (2)
with $K_{2}$ the modified Bessel function of the second kind. The parameter $\gamma_{\rm th}$ plays the role of an effective temperature $T$ of the quasi-thermal distribution ($\gamma_{\rm th}=kT/m_{e}c^{2}$, with $k$ the Boltzmann constant and $m_{e}$ the electron rest mass), while $\gamma_{\rm nth}$ in Eq.1 is the minimum Lorentz factor of the non-thermal tail, described as a power law with slope $p$ with cut-off at $\gamma_{\rm c}$ (with $\gamma_{\rm c}\gg\gamma_{\rm nth}$). These “injected” particles then cool and radiate.
@@
where $V$ is the volume of the emitting region. It is useful to introduce the fraction of energy contained in the non-thermal tail with respect to the total one:
@@
Figure 1: Electron energy distribution in the emission region calculated at different times (given in units of the light crossing time, $R/c$) with the parameters used for model A. The equilibrium distribution comprises a cooled low energy tail $N(\gamma)\propto\gamma^{-2}$, the high-energy tail (rapidly decreasing as a function of energy) of the thermal Maxwellian (which has the peak at $\gamma\simeq 100$) and the high-energy non-thermal power law with slope $p+1$ (with $p=2.3$ is the slope of the injected power law) up to the cut-off at $\gamma_{c}$.
@@ 2.1 Fixing parameters
Since we intend to explore the effects of the complex EED on the spectral energy distribution (SED) of FSRQ, we assume benchmark values for the physical parameters of the blazar emission region inspired by modeling of FSRQ (e.g. Tavecchio et al. 2000; Ghisellini et al. 2010). We assume $R=5\times 10^{16}$ cm, $B=1.6$ G, $\Gamma=20$, $\theta_{\rm v}=3.7$ deg (which combine to give a Doppler factor $D=15$). For the external radiation we assume that the dominant radiation field is that associated to the BLR, approximated (in the source rest frame) as a black body peaking at $\nu_{\rm ext}=2\times 10^{15}$ Hz (see the discussion in Tavecchio & Ghisellini 2008) with energy density $U_{\rm ext}=2.5\times 10^{-2}$ erg cm-3 (e.g. Ghisellini & Tavecchio 2008). We further assume that the source is located at the typical redshift $z=2$.
@@
A possible scenario for the origin of the shock at which electrons are heated and subsequently accelerated to high-energies is that invoking the interaction of different portions of the flow characterized by different speed (internal shocks, e.g. Spada et al. 2001). In this case we expect a mildly relativistic shock, with $\Gamma_{\rm u}\sim 1.5-2$. These shocks can efficiently accelerate particles only for low magnetization of the plasma (Sironi et al. 2015b). Indeed, FSRQ jets at the distance where the observed radiation is produced is expected to have a small magnetization (e.g. Sikora et al. 2005; Celotti & Ghisellini 2008). In these conditions we can rely on the results of Crumley et al. (2019), that report the analysis of PIC simulations for subluminal, high-Mach number, mildly relativistic ($\Gamma_{\rm u}=1.7$) weakly magnetized jets. The simulations show the development of a thermal component with average Lorentz factor around $\langle\gamma\rangle\approx 300$, and of a non-thermal power law with slope 2.2-2.3 containing a small fraction of the shock energy, $\epsilon_{e}\approx 10^{-3}$. However this last result could depend on the detailed configuration (i.e. inclination of the magnetic field and magnetization). PIC simulations of weakly magnetized, relativistic ($\Gamma_{\rm u}>3$) shocks show that the fraction of energy conveyed to non-thermal electrons can be much larger than that estimated in the mildly relativistic case, of the order of $\epsilon_{e}\approx 10^{-1}$ (Sironi & Spitkovsky 2011). These conditions could apply to FSRQ if the acceleration occurs at a stationary oblique recollimation shock (e.g. Bodo & Tavecchio 2018; Zech & Lemoine 2021).
@@ 3 Results
333 The last row reports $\epsilon_{e}$ derived from the input parameters with Eq. 7 and $\Gamma_{\rm u}=1.5$. The injected electron luminosity $L_{e}$ is in units of $10^{44}$ erg s-1.
@@ 4 Discussion
In this paper we have studied the case of the most powerful FSRQs, for which the thermal component naturally falls in the MeV band. For other types of blazars the situation is less straightforward. Less powerful FSRQ display the EC peak at higher energies (100 MeV-1 GeV, e.g. Marcotulli et al. 2022). If related to the thermal bump, such high peak energies would imply correspondingly high electron temperature, possibly related to larger dissipation parameter $\epsilon_{e}$. We remark that in this case, the high-energy part of the EC component would be affected by KN effects that smooth the spectrum, making difficult to identify the spectral structure (this effect is already visible in the D model in Fig. 2). At even lower power, blazars of the BL Lac type (LSP and HSP), display peaks above 10 GeV, which require electrons with $\gamma\gtrsim 10^{4}$ (e.g., Tavecchio et al. 2010), incompatible with thermal components from mildly relativistic shocks. In this case the thermal bump would instead appear in the low energy part of the IC component. We however remark that these sources produce high-energy radiation mainly through SSC, resulting in quite smooth spectra even for prominent thermal bumps (as also visible in Fig. 2).